OpenStudy (cinnamonbun):

How do I solve x-6y<_12 (<_ means less than or greater to) ~Graphing Inequalities~

OpenStudy (rayep12):

what do u think it is

OpenStudy (cinnamonbun):

I'm not sure, that's why I asked :c

OpenStudy (rayep12):

@TheSmartOne

jhonyy9 (jhonyy9):

x-6y <= 12 so this mean less or equal 12 than you need graphing this inequality so write it hence 6y >= x-12 y >= (x-12)/6 so for x = 0 get y = (-12)/6 = - 2 for x=12 will get y=0  hope this helped

jhonyy9 (jhonyy9):

@TheSmartOne opinion please - ty

OpenStudy (phi):

you could change the relation (as if it were an equation) except *no divide or multiply by a negative number* \[ x-6y\le12 \] solving for y: add 6 y to both sides \[ x \le 6y +12\] add -12 to both sides \[ x -12 \le 6y\] divide both sides by 6 \[ \frac{1}{6}x -2 \le y \\ y \ge \frac{1}{6}x -2\]

OpenStudy (phi):

you could plot this as the line \[ y = \frac{1}{6}x -2 \] and then, because it is really \[ y \ge \frac{1}{6}x -2 \] "shade in" the region above the line where y is "above" or bigger than the line

OpenStudy (-studywith-):

How do I solve x-6y<_12 (<_ means less than or greater to) ~Graphing Inequalities~ First, your slope intercept form: y = mx + b where m is the slope and m is the y-intercept. Try to put the equation x - 6y <= 12 in that form. Remember you switch the signs when multiplying/dividing by a negative number. So you have y => -2 + x/6 or y => x/6 -2. Now we have to graph that. Well, it's fairly simple now. 1/6 is your slope since mx = x/6 and if you divide x from that: x/6 * 1/x the x's cancel eachother out and you're left with 1/6. So this is the equation so far: y => 1/6 + b. And your b is just -2. So y => 1/6 - 2. Graph that. Need more help?